OpenStudy (anonymous):

find the solution set of x^2+3x+9=0

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

can you help me i missed a bunch of school and that is what we learned when i missed

terenzreignz (terenzreignz):

Surely :) If you have a quadratic equation of the form \[\large \color{red}ax^2 +\color{blue}bx+\color{green}c=0\] Then the solutions would be given by this formula~ \[\Large x=\frac{-\color{blue}b\pm\sqrt{\color{blue}b^2-4\color{red}a\color{green}c}}{2\color{red}a}\]

terenzreignz (terenzreignz):

Just identify correctly your values for a,b, and c, and plug them in :)

OpenStudy (anonymous):

ok... :/ is a 3 b 9 and c0

terenzreignz (terenzreignz):

No... look again~ a is the coefficient (number beside) \(x^2\) b is the coefficient of \(x\) c is the number with no x attached to it.

OpenStudy (anonymous):

a=1 b=3 and c=9

terenzreignz (terenzreignz):

Good :) Now just plug them into this formula. It would give you the two solutions you need :) \[\Large x=\frac{-\color{blue}b\pm\sqrt{\color{blue}b^2-4\color{red}a\color{green}c}}{2\color{red}a}\]

OpenStudy (anonymous):

\[-3\pm \sqrt{3^2-4(1)(9)}\]

OpenStudy (anonymous):

for the top row

terenzreignz (terenzreignz):

Very nice :) For the bottom "row" it's just 2(1) = 2, aye? ^.^

OpenStudy (anonymous):

not really but i got -27/2

terenzreignz (terenzreignz):

Okay... let's keep things real... why don't you simplify the inside of that radical first?

OpenStudy (anonymous):

is the answer -3+3isquareroot3/2, -3-3isquareroot3/2

terenzreignz (terenzreignz):

Yeah, that's it. I don't know how you got that, but who can deny its correctness? :P \[\Large x = \frac{-3\pm\sqrt{3^2-4(1)(9)}}{2(1)}=\frac{-3\pm \sqrt{-27}}{2}\] \[\Large x = \frac{-3\pm\sqrt{-1\cdot9\cdot3}}{2}=\frac{-3\pm3\color{red}i\sqrt{3}}{2}\]

OpenStudy (anonymous):

oh thank heavens hahaha